In a period of 1.50 s, 5.00 *10^23 nitrogen molecules strike a wall of area 8.20 cm2. If the molecules move at 328 m/s and strike the wall head on in a perfectly elastic collision, find the pressure exerted on the wall. (The mass of one N2 molecule is 4.68 10-26 kg.)
Answer #1

answer) we know the formula for impulse generated

F=ndP/dT=n*m*dv/t.........1)

since the collision is elastic

dv=328-(-328)=656m/s

so using eqn 1 we have

F=5*1023*4.68*10-26kg*656m/s)/1.5=10.2336N

we know the formula that

P=F/A=10.2336/8.20*10-4m2=12480 N/m2

so the answer is 12480 N/m2 or 1.25*104 N/m2 or Pa or 12kPa or 12.5 kPa or 1.2*104 N/m2 or Pa ( every answer is correct, its just depends how they want, if nothing is given just write the first answer or the second answer and Pa=N/m2)